OpenStudy (anonymous):

Show that there does not exist any number C such that the following function would be a probability mass function (p.m.f) f(x)=C/x, for x=1,2,3,4,... 0 elsewhere...

OpenStudy (anonymous):

did you try taking \(\sum_{n=1}^{\infty}\frac{C}{n}\)?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

which is the same as \[C\sum_{n=1}^{\infty}\frac{1}{n}\]

OpenStudy (anonymous):

series does not converge, so it cannot be 1

OpenStudy (anonymous):

yw